OpenStudy (anonymous):

If the circumference of the circle below is 54 cm, what is the length of arc XY?

OpenStudy (anonymous):



OpenStudy (amistre64):

what part of 360 do you have?

OpenStudy (anonymous):

WHAT

OpenStudy (amistre64):

\[54cm=360^o\] \[\frac{1}{360}54cm=1^o\] \[\frac{120}{360}54cm=120^o\]

OpenStudy (amistre64):

what part of 360 do they give you?

OpenStudy (amistre64):

you are told that you have 120 out of 360 degrees; this gives you a ratio of 120/360 to work with. since the full 360 degrees has a length of 54 then 120/360th of that is the answer

OpenStudy (anonymous):

so the answer is 18???

OpenStudy (amistre64):

120/360 = 1/3 54/3 is ... 1 drop a 2 and carry down the 4 ... 8 yes